# Automated Essay Scoring Incorporating Annotations from Automated Feedback Systems

### v1

@@ 3.2 Scoring Accuracy
Given the classification head is typically randomly initialized, we were also interested in whether these results were stable. We trained the scorer 10 times and reported the average, minimum, and maximum agreement levels for each agreement statistic we listed above. These statistics can be found in Table 6.
@@
All the models performed well above the human baseline. Out of the 10 separate trials, no model trained on the full-text alone scored as accurately as any of the models trained on the component annotated data or the data with both argumentative components and language errors annotated. An interesting observation is that the SMD, as calculated by (8), is only positive for models trained on the combined annotations and that the models trained on component annotated text showed the most controlled SMDs.
@@ 3.3 Potential bias
While these bias results were higher than expected, however, a cursory look seems to suggest that the use of combined annotations is mitigating some of the biases rather than exacerbating them.
This work is one of a number of works that highlight the growing need to address the bias introduced by automation, especially ELL students Ormerod et al. (2022). This suggests we need to apply bias mitigation, which could be of the form of a regression-based system with adjusted cut-off points Ormerod (2022b), or some sort of reinforcement learning mechanism.
@@ 4 Discussion
The results of this study demonstrate that incorporating feedback-oriented annotations into automated essay scoring (AES) pipelines can significantly improve scoring accuracy and provide meaningful, interpretable insights for students. The work of Uto and Uchida (2020) suggests this is also true for traditional global features. What we propose a realignment of AES to incorporate AWE elements so that we can provide students with more than just a score.

### v2

@@ 3.2 Scoring Accuracy
Given that the classification head is typically randomly initialized, we were also interested in whether these results were stable. We trained the scorer 10 times and reported the average, minimum, and maximum agreement levels for each agreement statistic we listed above. These statistics can be found in Table 6.
@@
All the models performed well above the human baseline. Out of the 10 separate trials, no model trained on the full-text alone scored as accurately as any of the models trained on the component annotated data or the data with both argumentative components and language errors annotated. An interesting observation is that the SMD, as calculated by (8), is only positive for models trained on the combined annotations, and that the models trained on component annotated text showed the most controlled SMDs.
@@ 3.3 Potential bias
While the resulting bias was higher than expected, a cursory look seems to suggest that the use of combined annotations is mitigating some of the biases rather than exacerbating them.
This work is one of a number of works that highlight the growing need to address the bias introduced by automation, especially for ELL students Ormerod et al. (2022). This suggests we need to apply bias mitigation, which could be of the form of a regression-based system with adjusted cut-off points Ormerod (2022b), or some sort of reinforcement learning mechanism.
@@ 4 Discussion
The results of this study demonstrate that incorporating feedback-oriented annotations into automated essay scoring (AES) pipelines can significantly improve scoring accuracy and provide meaningful, interpretable insights for students. The work of Uto and Uchida (2020) suggests this is also true for traditional global features. What we propose is a realignment of AES to incorporate AWE elements so that we can provide students with more than just a score.